OpenStudy (anonymous):

Solve the equation by factoring: x2 - 7x + 12 = 0

OpenStudy (anonymous):

x^2 - 4x - 3x + 12 = 0 x(x - 4) -3 (x - 4) = 0 (x - 3) (x - 4) = 0 x - 3 = 0 => x = 3 x - 4 = 0 => x = 4